# Accelerating Natural Gradient Descent for PINNs with Randomized Numerical Linear Algebra

### v2

@@ 2.2 Natural Gradient Descent
where $\mu>0$ is a regularization parameter, typically adapted during the optimization process.
2.3 Randomized Nyström preconditioning
The Gramian matrix $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ (2) is positive semidefinite and often numerically rank-deficient, as shown in Figure 1. In this setting, Randomized Nyström preconditioning [26] enables efficient construction of a preconditioner for $(\mathbf{G}+\mu\mathbf{I})$ using only matrix-vector products with $\mathbf{G}$. This preconditioner will be used with conjugate gradient to compute the regularized natural gradient (3).
Randomized Nyström approximation
@@ 2.3 Randomized Nyström preconditioning
Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be symmetric positive semidefinite (PSD). The randomized Nyström approximation [51, Section 14] with a random test matrix $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ is
@@
The randomness in $\boldsymbol{\Omega}$ ensures that $\hat{\mathbf{G}}_{\mathrm{nys}}$ approximates $\mathbf{G}$ well with high probability [51, Section 14.4]. In this work, we take $\boldsymbol{\Omega}$ to be standard normal. This choice, over sampling-based methods [2, 5] and other alternatives [4, 51, 53, 31], is motivated by stronger theoretical guarantees [26, 51] and by the fact that in our setting $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ in (11) is accessible only via matrix-vector products, and therefore structured test matrices requiring entrywise access are not applicable or less advantageous.
We emphasize that formula (4) is numerically unstable and should not be used directly. A more stable and efficient implementation, which also returns the eigendecomposition of $\hat{\mathbf{G}}_{\mathrm{nys}}$, is described in [51, Algorithm 16] and included in Appendix A.
Randomized Nyström preconditioner
Given an eigendecomposition $\hat{\mathbf{G}}_{\mathrm{nys}}=\mathbf{U}\hat{\boldsymbol{\Lambda}}\mathbf{U}^{\top}$ with $\hat{\boldsymbol{\Lambda}}=\mathrm{diag}(\hat{\lambda}_{1},\dots,\hat{\lambda}_{\ell})$ and $\hat{\lambda}_{1}\geq\dots\geq\hat{\lambda}_{\ell}$, the randomized Nyström preconditioner [26] for $(\mathbf{G}+\mu\mathbf{I})$ with $\mu\geq 0$ is
@@
To ensure good performance in expectation, the sampling parameter $\ell$ should be proportional to the effective dimension $d_{\mathrm{eff}}(\mu)=\sum_{i=1}^{p}\lambda_{i}(\mathbf{G})/(\lambda_{i}(\mathbf{G})+\mu)$, which acts as a smoothed count of eigenvalues of $\mathbf{G}$ larger than $\mu$ [26, Theorem 1.1]. When $\mathbf{G}$ exhibits strong spectral decay, $d_{\mathrm{eff}}(\mu)\ll p$, making the randomized Nyström preconditioner particularly effective.
@@
3 Matrix-free Natural Gradient Descent
Prior work has shown that matrix-vector products with the Gramian $\mathbf{G}(\boldsymbol{\theta})$ (2) can be computed via automatic differentiation for specific metrics [38, 55, 73]. We present a unified framework for metric selection and show that autodiff-based matvecs extend to broader metric classes.
3.1 Choice of the Metric
In some cases, the problem’s structure naturally suggests a metric, as in Variational MonteCarlo [58, 36, 55] or Sobolev gradient flows [35, 17]. In general, the link between NGD and Riemannian gradient descent [54] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian [14]. Further details are deferred to Appendix B.
Linear problems
Consider a linear problem $\mathcal{A}u=f\in\mathrm{W}^{\prime}$, where $\mathrm{V},\mathrm{W}$ are Hilbert spaces, and $\mathcal{A}:\mathrm{V}\to\mathrm{W}^{\prime}$ is a linear isomorphism. The associated least-squares problem and natural metric are
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\|{\mathcal{A}u-f}\right\|_{\mathrm{W}^{\prime}}^{2},\qquad{g}\left({v},\,{w}\right)\coloneqq\left({\mathcal{A}v},\,{\mathcal{A}w}\right)_{\mathrm{W}^{\prime}}.$ (6)
If $\mathrm{V}=\mathrm{W}$ and $\mathcal{A}$ is symmetric positive definite, one can endow $\mathrm{V}$ with the $\mathcal{A}$-induced inner product $\left({v},\,{w}\right)_{\mathcal{A}}=\left\langle{\mathcal{A}v},\,{w}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$, and use it as the NGD metric. Then, up to an additive constant irrelevant for optimization, the loss (6) coincides with the variational form used in the Deep Ritz method [25]
$\inf_{u\in\mathrm{V}}E(u)\coloneqq\frac{1}{2}\left\langle{\mathcal{A}u},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}-\left\langle{f},\,{u}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}.$ (7)
A link with operator preconditioning
NGD can be viewed as a preconditioned gradient descent, where the preconditioner is the Gramian [21]. For a linear parametrized model $u_{\boldsymbol{\theta}}=\sum_{i=1}^{p}\theta_{i}\,\varphi_{i}$, the Gramian becomes $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\varphi_{i}},\,{\varphi_{j}}\right)$. If $(\mathcal{M},g)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ and $g$ is induced by a positive definite linear operator $\mathcal{G}:\mathrm{V}\to\mathrm{V}^{\prime}$, then $\mathbf{G}(\boldsymbol{\theta})=\mathbf{G}$ where $[\mathbf{G}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ does not depend on $\boldsymbol{\theta}$. This coincides with preconditioners arising from an operator preconditioning perspective, in the context of Galerkin discretizations of PDEs [37, 48, 43, 30].
Nonlinear problems
If $\mathcal{M}$ is equipped with a metric $a$ and $\mathrm{Hess}^{a}E(u)$ is positive definite for all $u\in\mathcal{M}$, Newton’s method [14] is induced by choosing the metric
${g_{u}}\left({v},\,{w}\right)=\mathrm{Hess}^{a}E(u)[v,w].$ (8)
In the case $(\mathcal{M},a)=(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ for a Hilbert space $\mathrm{V}$, this leads to the Energy Natural Gradient Descent (ENGD) method [54, 55]. Note that to interpret ENGD as a Natural Gradient Descent, it is necessary that $\mathrm{Hess}^{a}E(u)$ is SPD for all $u\in\mathcal{M}$, as otherwise (8) does not define a valid metric.
Consider now the nonlinear problem $\mathcal{R}(u)=0$, where $\mathcal{R}:\mathrm{V}\to\mathrm{W}$ is a (possibly nonlinear) residual between Hilbert spaces $\mathrm{V},\mathrm{W}$. For the least-squares functional $E(u)=\frac{1}{2}\left\|{\mathcal{R}(u)}\right\|_{\mathrm{W}}^{2}$, a symmetric positive semidefinite approximation of the Hessian is the Gauss-Newton one [1, 70]:
This choice of the metric leads the Gauss-Newton NGD method [38, 33]. Note that $g_{u}$ is, in general, only semidefinite. Positive definitness of $g_{u}$ is necessary to interpret it as a Riemannian metric, and whether this condition holds depends on the structure of the underlying nonlinear PDE.
3.2 Matrix-free formulation
We show that matrix-vector products with the Gramian (2) can be computed via automatic differentiation, provided the metric satisfies the following assumption.
Assumption 3.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and let $\mathcal{M}\subseteq\mathrm{V}$, where $\mathrm{V}$ is a Hilbert space of functions on $\Omega$. Assume that the metric $g_{u}$ takes the form
for all $v,w\in\mathrm{T}_{u}\mathcal{M}\subseteq\mathrm{V}$, where $\mathcal{F}_{u}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi)^{k}$ is linear and continuous.
Theorem 3.2.
Let 3.1 hold, and let $\mathtt{sg}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ denote the stop-gradient operator, formally defined by $\mathtt{sg}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\nabla\mathtt{sg}(\boldsymbol{\theta})=\mathbf{0}$. If the integral (10) is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
$\begin{split}\mathbf{G}(\boldsymbol{\theta})&=\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta})^{\top}\begin{bmatrix}w_{1}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{1})\mathbf{I}_{k}&&\\ &\ddots&\\ &&w_{q}\varphi_{u_{\boldsymbol{\theta}}}(\mathbf{x}_{q})\mathbf{I}_{k}\end{bmatrix}\mathbf{J}_{\boldsymbol{\theta}}\mathbf{F}(\boldsymbol{\theta}),\quad\,\,\,\mathbf{F}(\boldsymbol{\theta})=\begin{bmatrix}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{1})\\ \vdots\\ \mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}(\mathbf{x}_{q})\end{bmatrix},\end{split}$ (11)
The proof of 3.2 is deferred to Appendix C. As a consequence of 3.2, matrix-vector products with $\mathbf{G}(\boldsymbol{\theta})$ can be computed using only products with the Jacobian and the transpose Jacobian of $\mathbf{F}$. These operations can be efficiently implemented via automatic differentiation, using forward-mode autodiff for Jacobian-vector products (JVPs) and backward-mode autodiff for vector-Jacobian products (VJPs). Note that most autodiff frameworks include a stop-gradient function.
Remark 3.3.
3.1 encompasses standard metrics used in PDE-related function spaces. For instance, the $\mathrm{L}^{2}(\Omega)$ and $\mathrm{H}^{1}(\Omega)$ metrics correspond to the choices $\mathcal{F}_{u}v=v$ and $\mathcal{F}_{u}v=(\nabla v,v)$, respectively. It also includes the Gauss-Newton metric (9) on $\mathrm{L}^{2}(\Omega)$, obtained by taking $\mathcal{F}_{u}=\mathrm{D}\mathcal{R}(u)$, for which matrix-free NGD implementations have been proposed in [38, 55] for specific problems. However, 3.1 is significantly more general and covers a much broader class of metrics. In practice, it includes many of the metrics relevant in the context of PDEs, since these typically need to be expressible as integrals to be computationally tractable. For example, it includes the Newton metric ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+3u^{2}vw$ in [55, Section 4.2], and the energy inner product for the Gross–Pitaevskii equation as considered in [35], ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+Vvw+\beta\lvert u\rvert^{2}vw$.
Computational Cost
With automatic differentiation, the cost of each JVP and VJP is a small constant factor (typically between $2$ and $4$) times the cost of evaluating $\mathbf{F}_{i}(\boldsymbol{\theta})$ for $i=1,2$ [29, 49]. For a multilayer perceptron, the evaluation cost scales as $\mathcal{O}(pq)$, where $p$ is the number of trainable parameters and $q$ is the number of quadrature points. Therefore, the cost of each matrix-vector product with $\mathbf{G}(\boldsymbol{\theta})$ is $\mathrm{C}_{\mathrm{mv}}=\mathcal{O}(pq)$. Note however that when executed on modern GPUs, batch evaluation across $q$ points benefits from parallelization.

### v3

@@ 3.1 Natural Gradient Descent
where $\mu>0$ is a regularization parameter, typically adapted during the optimization process. Once the (regularized) natural gradient descent is computed, the parameters are updated as
$\boldsymbol{\theta}\leftarrow\boldsymbol{\theta}-\alpha\mathbf{d}=\boldsymbol{\theta}-\alpha(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})^{-1}\nabla_{\boldsymbol{\theta}}L(\boldsymbol{\theta}),$
where the step-size $\alpha$ is typically chosen via a line-search procedure.
3.2 A matrix-free formulation
Leaving aside the resolution of the least squares problem, even assembling the Gramian $\mathbf{G}(\boldsymbol{\theta})$ in (5) can be computationally prohibitive in both flops and memory. Indeed, $\mathbf{G}(\boldsymbol{\theta})$ is a $p\times p$ matrix, where $p$ is the number of network parameters, which can be in the order tens of thousands even for relatively small neural networks. To mitigate this issue, we consider a matrix-free approach: $\mathbf{G}(\boldsymbol{\theta})$ is never explicitly assembled, and only matrix-vector products are computed as needed. This is particularly efficient when solving linear systems with $(\mathbf{G}(\boldsymbol{\theta})+\mu\mathbf{I})$ using iterative methods such as Conjugate Gradient (CG), which require only matvecs.
Prior work has demonstrated that matrix-vector products with the Gramian (5) can be computed via automatic differentiation for specific metrics [46, 65, 86]. Here, we show that autodiff-based matvecs extend to broader metric classes. Specifically, we show that matrix-vector products with the Gramian (5) can be computed via automatic differentiation, provided the metric satisfies the following assumption and it is discretized via quadrature.
Assumption 3.1.
Let $\Omega\subset\mathbb{R}^{d}$ be a domain, and let $\mathrm{V}$ be a Hilbert space of functions on $\Omega$. Assume that the metric $g_{u}$ takes the form
for all $v,w\in\mathrm{T}_{u}\mathrm{V}\subseteq\mathrm{V}$, where $\mathcal{F}_{u}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi_{u})^{k}$ is linear and continuous.
Theorem 3.2.
Let ˜3.1 hold, and let $\mathtt{sg}:\mathbb{R}^{p}\to\mathbb{R}^{p}$ denote the stop-gradient operator, formally defined by $\mathtt{sg}(\boldsymbol{\theta})=\boldsymbol{\theta}$ and $\nabla\mathtt{sg}(\boldsymbol{\theta})=\mathbf{0}$. If the integral (7) is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
[…]
Proof.
Let $\mathbf{e}_{h}\in\mathbb{R}^{p}$ denote the $h$-th vector of the canonical basis, i.e., the vector that has a $1$ in the $h$-th entry and 0 in all other entries. Using the linearity and continuity of $\mathcal{F}_{u_{\boldsymbol{\theta}}}$,
or, in matrix form,
where the Hilbert adjoint $\left(\mathrm{D}_{\boldsymbol{\theta}}\mathcal{F}_{u_{\mathtt{sg}(\boldsymbol{\theta})}}u_{\boldsymbol{\theta}}\right)^{*}$ is taken with respect to the inner product on $\mathrm{L}^{2}(\Omega,\varphi_{u})^{k}$, and we identify a matrix with the induced linear map.
and thus, in matrix form:
As a consequence of ˜3.2, matrix-vector products with $\mathbf{G}(\boldsymbol{\theta})$ can be computed using only products with the Jacobian and the transpose Jacobian of $\mathbf{F}$. These operations can be efficiently implemented via automatic differentiation, using forward-mode autodiff for Jacobian-vector products (JVPs) and backward-mode autodiff for vector-Jacobian products (VJPs). Note that most autodiff frameworks include a stop-gradient function.
Computational Cost
With automatic differentiation, the cost of each JVP and VJP is a small constant factor times the cost of evaluating $\mathbf{F}(\boldsymbol{\theta})$ [34, 58]. For a feedforward neural network, the evaluation cost scales as $\mathcal{O}(pq)$, where $p$ is the number of trainable parameters and $q$ is the number of quadrature points. Therefore, the cost of each matrix-vector product with $\mathbf{G}(\boldsymbol{\theta})$ is $\mathrm{C}_{\mathrm{mv}}=\mathcal{O}(pq)$. Note however that when executed on modern GPUs batch evaluation across $q$ points benefits from parallelization.
Remark 3.3.
More in general, it can be shown that if the metric $g_{u}$ takes the form
where $\mathcal{F}_{u}^{i}:\mathrm{V}\to\mathrm{L}^{2}(\Omega;\varphi_{u})^{k}$ are linear and continuous for $i=1,2$, and the integral is approximated using a quadrature rule with points $\{\mathbf{x}_{r}\}_{r=1}^{q}$ and weights $\{w_{r}\}_{r=1}^{q}$, then
[…]
Remark 3.4.
˜3.1 encompasses standard metrics used in PDE-related function spaces. For instance, the $\mathrm{L}^{2}(\Omega)$ and $\mathrm{H}^{1}(\Omega)$ metrics correspond to the choices $\mathcal{F}_{u}v=v$ and $\mathcal{F}_{u}v=(\nabla v,v)$, respectively. It also includes the Gauss-Newton metric on $\mathrm{L}^{2}(\Omega)$, for which matrix-free NGD implementations have been proposed in [46, 65] for specific problems. However, ˜3.1 is significantly more general and covers a much broader class of metrics. In practice, it includes many of the metrics relevant in the context of PDEs, since these typically need to be expressible as integrals to be computationally tractable. For example, it includes the Newton metric ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+3u^{2}vw$ in [65, Section 4.2], and the energy inner product for the Gross–Pitaevskii equation as considered in [41], ${g_{u}}\left({v},\,{w}\right)=\int_{\Omega}\nabla v\cdot\nabla w+Vvw+\beta\lvert u\rvert^{2}vw$.
3.2.1 A link with Operator Preconditioning and the case of Finite Elements Interpolated Neural Networks
The Gramian (5) can be interpreted in the context of operator preconditioning [44, 57, 52, 36]. Specifically, NGD can be seen as a preconditioned gradient descent method, with the Gramian serving as the preconditioner [24]. For a linear parametrized model $u_{\boldsymbol{\theta}}=\sum_{i=1}^{p}\theta_{i}\,\varphi_{i}$, the Gramian reduces to $\left[\mathbf{G}(\boldsymbol{\theta})\right]_{ij}={g_{u_{\boldsymbol{\theta}}}}\left({\varphi_{i}},\,{\varphi_{j}}\right)$. If $(\mathrm{V},\left({\cdot},\,{\cdot}\right)_{\mathrm{V}})$ is a Hilbert space and $g$ is induced by a positive definite linear operator $\mathcal{G}:\mathrm{V}\to\mathrm{V}^{\prime}$, then
$\mathbf{G}(\boldsymbol{\theta})]_{ij}=[\mathbf{G}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}},$
which is independent of $\boldsymbol{\theta}$. This coincides exactly with preconditioners arising from operator preconditioning in Galerkin discretizations of PDEs [44].
The connection with operator preconditioning highlights that the choice of metric crucially influences the convergence speed of NGD. In general, the link between NGD and Riemannian gradient descent [64] suggests selecting the metric $g$ to reduce the condition number of the Riemannian Hessian. Further details are provided in Appendix˜A, where we present a unified framework for metric selection.
In the case of FEINNs, ˜3.2 can be specialized to connect with operator preconditioning for finite elements. Assume, for simplicity, that the interior nodes of the mesh are the first $N_{I}$, so that the FEINN can be written as $u_{\boldsymbol{\theta}}=\bar{u}_{h}+\sum_{i=1}^{N_{I}}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{i})\varphi_{i}$. Then the Gramian (5) takes the form
$\mathbf{G}(\boldsymbol{\theta})=\left(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{P}(\boldsymbol{\theta})\right)^{\top}\mathbf{G}_{h}\left(\mathbf{J}_{\boldsymbol{\theta}}\mathbf{P}(\boldsymbol{\theta})\right),\qquad\text{where}\qquad\mathbf{P}(\boldsymbol{\theta})=\begin{bmatrix}u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{1})\\ \vdots\\ u_{\boldsymbol{\theta}}^{\mathcal{NN}}(\mathbf{x}_{N_{I}})\end{bmatrix}$,
where $\mathbf{G}_{h}\in\mathbb{R}^{N_{I}\times N_{I}}$ with $\mathbf{G}_{h}]_{ij}=\left\langle{\mathcal{G}\varphi_{i}},\,{\varphi_{j}}\right\rangle_{\mathrm{V}^{\prime}\times\mathrm{V}}$ is the FEM operator preconditioner. Note that computing matvecs with the gramian requires only matvecs with $\mathbf{G}_{h}$, but not with its inverse, and JVPs and VJPs with $\mathbf{P}(\boldsymbol{\theta})$, which is an evaluation of the neural network at the interior nodes of the mesh.
Given that matrix-vector products with the Gramian $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ can be computed efficiently, it is appealing to compute the regularized natural gradient (6) by solving the linear system with an iterative method. Since $\mathbf{G}+\mu\mathbf{I}$ is symmetric positive definite, the method of choice is Conjugate Gradient (CG). In practice, $\mathbf{G}$ is often ill-conditioned, causing slow convergence of CG unless a suitable preconditioner is employed. However, the matrix $\mathbf{G}$ lacks the sparsity or grid-based structure typical of FEM/FD matrices, which are crucial for classical preconditioners such as Incomplete Cholesky, multigrid/multilevel methods, and domain decomposition techniques. Instead, the main characteristics of $\mathbf{G}$ are that it is symmetric positive semidefinite, typically exhibits rapid singular value decay, and can only be accessed through matrix-vector products. In this setting, Randomized Nyström [77, 31] is, to the best of our knowledge, the most effective method currently available for constructing a preconditioner. In this section, we first review randomized Nyström preconditioning, then present our proposed method, NyströmNGD, which combines matrix-free NGD with randomized Nyström.
4.1 Randomized Nyström approximation and preconditioning
@@
Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be symmetric positive semidefinite (PSD). The randomized Nyström approximation [60, Section 14] with a random test matrix $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ is
@@
The randomness in $\boldsymbol{\Omega}$ ensures that $\hat{\mathbf{G}}_{\mathrm{nys}}$ approximates $\mathbf{G}$ well with [60, Section 14.4]. In this work, we take $\boldsymbol{\Omega}$ to be standard normal. Then the following guarantees hold.
Proposition 4.1.
[60, Theorem 14.1] Let $\mathbf{G}\in\mathbb{R}^{p\times p}$ be a symmetric positive semidefinite matrix with eigenvalues $\lambda_{1}\geq\lambda_{2}\geq\dots\geq\lambda_{p}\geq 0$. Let $\boldsymbol{\Omega}\in\mathbb{R}^{p\times\ell}$ be a standard Gaussian matrix with $\ell<p$. Let $\hat{\mathbf{G}}_{\mathrm{nys}}$ be the randomized Nyström approximation defined in (11), and let $\hat{\mathbf{G}}_{k}$ be its truncation to rank $k$. Then, for all $k<\ell-1$, the error in expectation satisfyes:
$\mathbb{E}\left[\|{\mathbf{G}-\hat{\mathbf{G}}_{k}}\|_{2}\right]\leq\lambda_{k+1}+\frac{k}{\ell-k-1}\left(\sum_{j>k}\lambda_{j}\right).$
Hence, the Nyström approximation on average achieves the optimal error $\lambda_{k+1}$ for a rank-$k$ approximation, up to an additive term that depends on the tail of the eigenvalues and decreases with $\ell$. This approximation is particularly effective when $\mathbf{G}$ has a strong spectral decay, as it is in our setting. We emphasize that formula (11) is numerically unstable and should not be used directly. A more stable and efficient implementation, which also returns the eigendecomposition of $\hat{\mathbf{G}}_{\mathrm{nys}}$, is described in [60, Algorithm 16] and reported in Algorithm˜1.
Remark 4.2.
The choice of a Gaussian test matrix $\boldsymbol{\Omega}$ over sampling-based methods [2, 6] and other alternatives [5, 60, 63, 37], is motivated by stronger theoretical guarantees [31, 60] and by the fact that in our setting $\mathbf{G}=\mathbf{G}(\boldsymbol{\theta})$ in (8) is accessible only via matrix-vector products, and therefore structured test matrices requiring entrywise access are not advantageous.
[…]
Algorithm 1 Randomized Nyström Approximation [31, 60]
[…]
@@
To ensure good performance in expectation, the sampling parameter $\ell$ should be proportional to the effective dimension $d_{\mathrm{eff}}(\mu)=\sum_{i=1}^{p}\lambda_{i}(\mathbf{G})/(\lambda_{i}(\mathbf{G})+\mu)$, which acts as a smoothed count of eigenvalues of $\mathbf{G}$ larger than $\mu$ [31, Theorem 1.1]. When $\mathbf{G}$ exhibits strong spectral decay, $d_{\mathrm{eff}}(\mu)\ll p$, making the randomized Nyström preconditioner particularly effective. The following proposition makes this precise.
Proposition 4.3.
[31, Theorem 5.1] Suppose that the randomized Nyström preconditioner (12) is constructed using Algorithm˜1 with a Gaussian test matrix and $\ell=2\left\lceil 1.5d_{\mathrm{eff}}(\mu)\right\rceil+1$. Then, using $\mathbf{P}^{-1}$ as a preconditioner for $(\mathbf{G}+\mu\mathbf{I})$, results in the condition number bound in expectation:
$\mathbb{E}\left[\kappa_{2}(\mathbf{P}^{-1/2}(\mathbf{G}+\mu\mathbf{I})\mathbf{P}^{-1/2})\right]<28.$
@@
4.2 NyströmNGD